OpenStudy (anonymous):

If 400 pa of pressure is applied to an area of 55 m^2, which is the resulting force

OpenStudy (anonymous):

Since Pressure = Force/Area, to find the force you take Pressure * Area. Then check your units. 1 Pa = 1N/M^2

OpenStudy (shamim):

Above solution is correct!

OpenStudy (shamim):

Same thing i m telling Use p=F/A F=pA=?

OpenStudy (anonymous):

Pressure = Force / area, which means Force = pressure x area. Force = 400×55 = 22000N